The 28th term of an AP is - 5.find it's common difference if it's first term is 31

Answer by Jeetbhatt10th(11)

a28=-5=a+(n-1)d,
where a is the first term = 31 and d is the common diff.
=>-5=31+(28-1)d,
=>-5=31+27d
=>27d=-5-31
=>27d=-36
=>d=-36/27
=>d=-4/3 (answer)